Just #2, 14, 16. Thank you
EXERCISE SET 1.2 Practice Exercises In Exercises 1-16, solve and check each linear equation. 1. 7x 5 72 3. 11x - (6x -5) 40 5x (2x 10) 35 6.3x+5=2x+13 8. 13x + 14 = 12x-5 7. 7x 4x + 16 9. 3(r 2) +7 2(x + 5) 10. 2(x 1) +3 - 3(r + 1) 11. 3(x-4)-4(x _ 3) = x + 3-(x-2) 12. 2-(7x +5) 13 3x 13. 16 3(x 1) - (r -7) 14.5x-(2x +2) = x + (3x-5) 15. 25 [2 +5y 3y 2)]- -3(2y - 5)- [5(v 1)-3y +3] 16.)45 - [4- 2y 4(y 7)]- -4(1 + 3y)- [4.-3(y + 2) 2(2y 5)] Exercises 17-30 contain linear equations with constants in
Answer #1

2 6x -3 63 6x-66 X=// リ 5%-(2x+2)=x+(3x-5) 3x-2 = 4x-5 +2 +2 Y+2)-2(2y-5 -69 69 8l